OpenStudy (anonymous):

(5x-y)^4 +4y^3 = 20704 solve for (2,-2) find the derivative

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy!!! :) I saw that you asked for help in the chat, maybe try adding the url link so that helpers can find it easier :) I'm not very good at calculus, but I'll do my best to help :) so I'm guessing you're confused by the fact that you're supposed to take the derivatives with ys involved?

OpenStudy (anonymous):

there are 2 x and y

jigglypuff314 (jigglypuff314):

the derivative of y is y' or dy/dx so d/dx (5x - y)^4 = (chain rule) -> 4(5x - y)^3 (5 - dy/dx) or 4(5x-y)^3 (5-y' )

OpenStudy (anonymous):

ah ok ... would it be then 4(5x-y)^ (5-y') + 12y^2

OpenStudy (anonymous):

plus the y' to the 12y^2*y'

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

then you just plug in the points to solve for it

jigglypuff314 (jigglypuff314):

yes :) and know that derivative of 20704 = 0

OpenStudy (anonymous):

im kind of confused at this part... bc when i plug it in it gives me a big number but doesn't make sense

jigglypuff314 (jigglypuff314):

hmmm 4(5(2)-(-2))^3 (5-y') + 12(-2)^2 y' = 0

jigglypuff314 (jigglypuff314):

4(5(2)-(-2))^3 (5-y') + 12(-2)^2 y' = 0 4(12)^3 (5-y') + 12(4) y' = 0 6921 (5-y') + 48 y' = 0 34560 - 6921 y' + 48 y' = 0

OpenStudy (anonymous):

yea for 4((5*2)-(-2)))^3 gives me = 6912

jigglypuff314 (jigglypuff314):

oh yeah whoops that's right, typo on my part sorry

OpenStudy (anonymous):

thats exactly wat i got ... is that my answer

jigglypuff314 (jigglypuff314):

you would solve for y'

OpenStudy (anonymous):

ok thanks i got it!

jigglypuff314 (jigglypuff314):

glad I could help :)

OpenStudy (anonymous):

maybe you could help me with this one ... im sorry lol would you do product rule or just go ahead a do the derivative 2x^2-3xy+3y^3

jigglypuff314 (jigglypuff314):

yeah... product rule for the -3xy

OpenStudy (anonymous):

with 3y^3 right?

jigglypuff314 (jigglypuff314):

huh?

OpenStudy (anonymous):

(-3y') * 3y^3 + (9y^2*y')* (-3xy) ?

jigglypuff314 (jigglypuff314):

ermm... I'm slightly confused... I got 4x - 3 - 3y' + 9y^2 y'

jigglypuff314 (jigglypuff314):

whoops I meant 4x - 3y - 3xy' + 9y^2 y'

OpenStudy (anonymous):

ah ok .... i was confused to wat im supposed to use the product rule to

OpenStudy (anonymous):

the derivative of -3xy is not -3*y'

jigglypuff314 (jigglypuff314):

product rule -3xy = -3 (x*y) derivative -> -3 (y + xy')

OpenStudy (anonymous):

thanks you're really helpful

jigglypuff314 (jigglypuff314):

you're welcome ^_^